OpenStudy (anonymous):

A cylinder and a cone have the same diameter: 8 inches. The height of the cylinder is 5 inches. The height of the cone is 15 inches. Use π = 3.14. What is the relationship between the volume of this cylinder and this cone? Explain your answer by determining the volume of each and comparing them. Show all your work.

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

Volume of a cylinder: \(V = \pi r^2 h\) \(V = (3.14)(4)^2(5)\) What's 4 * 4?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

Yes, and what's 3.14 * 16 * 5?

OpenStudy (anonymous):

252.8

OpenStudy (anonymous):

Check again

OpenStudy (anonymous):

251.5

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

It's actually 251.2 Volume of a cone: \(V = \dfrac{1}{3} \pi r^2 h\) \(V = \dfrac{1}{3} (3.14) (4)^2 (15)\) What's 1/3 * 3.14 * 16 * 15?

OpenStudy (anonymous):

251.2

OpenStudy (anonymous):

Yes, they are the same.

OpenStudy (anonymous):

So the relationship is that they're the same.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

And the volume of a cone is ONE THIRD the volume of a cylinder. That's what you can say

OpenStudy (anonymous):

@ShawnSpencer Got it..?